# Under-coverage in high-statistics counting experiments with finite MC samples

### v2

@@ Abstract
We consider the problem of setting confidence intervals on a parameter of interest from the maximum-likelihood fit of a physics model to a binned data set with a large number of bins, large event-counts per bin, and in the presence of systematic uncertainties modeled as nuisance parameters. We use the profile-likelihood ratio for statistical inference and focus on the case in which the model is determined from Monte Carlo simulated samples of finite size. We start by presenting a toy model in which the asymptotic properties of the profile-likelihood ratio are manifestly broken even if the numbers of events per bin in both the data and simulated samples are seemingly large enough to warrant their validity. We then move to the general setting to show how statistical uncertainties in the Monte Carlo predictions can affect the coverage of confidence intervals constructed in the asymptotic approximation always in the same direction, namely they lead to systematic under-coverage.
arxiv: 2401.10542
@@ 1 Introduction
Due to the complexity of modeling particle physics processes from generator up to detector-level, the function ${\bf f}(\mu,\boldsymbol{\theta})$ can rarely be computed analytically, leaving Monte Carlo (MC) techniques as the only viable way for its determination. Given that MC samples necessarily comprise a finite number of events, a MC-simulated model prediction will never be exact, thus introducing an additional source of systematic uncertainty. A general rule is that the size of the samples should exceed that of the data by at least a few times. However, such a request can turn out computationally expensive or even unfeasible for analyses performed on high-statistics samples of events, especially when CPU-intensive MC simulations are needed to achieve the desired level of theoretical accuracy. In such cases, one has to live with samples of MC-simulated events of size comparable or smaller than the data. The situation can become even more problematic when the simulated events are weighted Glusenkamp .
@@
The correct frequentist approach to deal with the finite size of MC samples was first introduced by Barlow and Beeston in ref. BARLOW1993219 . In their seminal paper, the authors proposed the idea of treating the uncertain MC predictions as additional nuisance parameters of which the available MC sample represents an ancillary measurement. Since then, several improvements in the treatment of MC uncertainties have been proposed in the literature, with particular emphasis on the case of low-statistic samples and the proper handling of event weights BARLOW1993219 ; chirkin2013likelihooddescriptioncomparingdata ; Glusenkamp ; Arguelles .
@@
In this paper, we focus on the problem of setting confidence intervals on a parameter of interest in the presence of nuisance parameters and in the case of high-statistics counting experiments, for which asymptotic formulas from the theory of maximum likelihood estimators Cowan ; Kendall ; Asimov are expected to hold. We notice that the high-statistics case has received less attention in the literature compared to the low-statistics limit, where asymptotic formulas are more likely invalidated by low event counts. Furthermore, the high-statistics case is presently of great interest for precision measurements performed on large-scale data samples, such as those collected at the LHC, see e.g. refs. ATLAS:2024wla ; CMS:2024lrd for a recent review on this subject.
@@
A relevant example of a precision measurement performed on a large-scale data sample, and with a large number of nuisance parameters, is represented by the measurement of the $W$ boson mass by the CMS Collaboration CMS:2024lrd : the CMS measurement analyzed a sample of LHC collision data consisting of about $10^{8}$ single-muon events, binned in the kinematics and charge of the muons, resulting in approximately $3\times 10^{3}$ bins, which are used to simultaneously measure the mass of the $W$ boson together with a few thousands of nuisance parameters via a maximum-likelihood fit. The model is constructed by using samples of MC-simulated events whose overall statistical power exceeds that of the data by a factor of about $4$. Alternate MC predictions corresponding to ±$\pm 1\sigma$ variations of each nuisance parameter are constructed by reweighting the nominal MC simulation. The effect of statistical fluctuations in the alternate templates used to construct the systematic variations was observed to lead to a moderate under-coverage of Hessian uncertainties, which required an inflation of the statistical uncertainty of the MC templates to restore the correct coverage CMS:2024lrd . The main scope of this paper is to study the origin of such an effect from a more general perspective.
@@
This paper is organized as follows. We start by presenting a hypothetical experiment which can be considered paradigmatic of our case of interest: the data set consists in a histogram with a large number of bins; the event counts in each bin are large; the physics model used to describe the data includes nuisance parameters and is determined from MC-simulated samples of size comparable to the data. We compute numerically the coverage of intervals constructed from CI-setting methods that rely on the asymptotic properties of maximum-likelihood estimators, such as those based on the Hessian matrix or the profile-likelihood ratio, and find that such intervals cover systematically less than their claimed confidence level. Then, we try alternative CI-setting methods that do not rely on asymptotic properties, such as those inspired by the unified approach by Feldman and Cousins FC and again find similar issues. Inspired by our findings, we propose a heuristic definition of a confidence interval that better approaches the correct Feldman-Cousins interval. In the second part of the paper, we study the same problem from a more general perspective to ascertain in which circumstances statistical fluctuations of the MC model can lead to a systematic under-coverage of asymptotic confidence intervals. We conclude by suggesting numerical tests to quantify the impact of MC statistical fluctuations for a generic problem, remarking that the assumption of being in the asymptotic regime may not be simply guaranteed by the large size of the data and simulated samples alone.
@@ 2 A paradigmatic toy model
Consider a hypothetical experiment where two processes, respectively labeled with index $j=1,2$, generate independent, Poisson-distributed event counts $y_{ji}$ with expectation values in the $i^{\rm th}$ bin given by
$\langle y_{1i}\rangle\equiv\nu_{1i}=K\;\;\;{\rm and}\;\;\;\langle y_{2i}\rangle\equiv\nu_{2i}=\begin{cases}K\cdot(1+\epsilon)&\;\;{\rm for}\;i<\frac{n}{2}\\ K&\;\;{\rm for}\;i\geq\frac{n}{2}\end{cases}$ (1)
where $|\epsilon|\ll 1$, $K\gg 1$, and $n$ is the (even) number of analyzed bins. In a real experiment the constant $K$ would be related to e.g. beam flux, detector efficiency, cross sections, branching ratios, etc. In our study, we find it more convenient to trade it for the total number of expected events $N$ in the data sample.
Let the purpose of this hypothetical experiment be a measurement of the total number of events from the process $j=1$, which can be thought of as the “signal”, with process $j=2$ playing the role of a “background” of unknown cross section. It is easy to see how the data could be reduced to just two bins with no loss of information on the parameter of interest. However, the crucial assumption made here is that the true model of eq. (1) is not known a priori but can only be predicted from finite-size MC samples consisting of ${t}_{ji}$ simulated events in the $i^{\rm th}$ bin and for the $j^{\rm th}$ process. Since $t_{ji}$ are also random variables, their values will differ from $\nu_{ij}$ by statistical fluctuations introduced by the MC generation, which could indeed justify preserving the full granularity of the data.
Given a set of MC predictions $t_{ji}$, the signal normalization can be extracted by parameterizing the expectation value $\langle y_{i}\rangle$ as:
@@
where strength modifiers $\mu_{j}$ have been introduced for the two processes as a proxy for the total number of expected events, and $k_{j}$ are constants that rescale the normalization of the MC samples to some reference value. Rescaled MC predictions $T_{ji}=t_{ji}/k_{j}$, commonly known as templates, provide the elements of the Jacobian matrix of ${\bf f}$ with respect to its parameters. We choose $k_{j}$ equal to the ratio between the expectation value of $t_{ji}$ and the true value $\nu_{ji}$, so that by construction $f_{i}$ is an unbiased estimator of $\langle y_{i}\rangle$ for $\mu_{1}\equiv\mu_{1,{\rm t}}=1$ and $\mu_{2}\equiv\mu_{{2,\rm t}}=1$. Consistently with this assumption, values for $y_{ji}$ and $t_{ji}$ can be generated by drawing random integer numbers from the Poisson density with mean values $\nu_{ji}$ and $k_{j}\nu_{ji}$, respectively, that is:
${\rm Prob}\left[y_{ij}\right]=\frac{e^{-\nu_{ji}}\left(\nu_{ji}\right)^{y_{ji}}}{y_{ij}!}\;\;\;{\rm and}\;\;\;{\rm Prob}\left[t_{ij}\right]=\frac{e^{-k_{j}\nu_{ji}}\left(k_{j}\nu_{ji}\right)^{t_{ji}}}{t_{ij}!}$ (3)
Examples of templates $T_{ji}$ are shown in figure 1 for representative values of $N$, $\epsilon$, $n$, and $k_{j}$.
Figure 1: Examples of templates $T_{ji}$ generated as described in the text for $N=2\times 10^{6}$, $\epsilon=0.03$, $n=200$, and for two values of $k_{1}=k_{2}=k$, namely $k=1$ (left) and $k=10$ (right).
@@
The coverage of $I$ is defined as the fraction of pseudo-experiments for which $\mu_{\rm t}\in I$. In order to verify that the coverage of $I$ does not depend on $\mu_{\rm t}$, two additional independent sets of data ${\bf y}_{\pm}$ are considered. The latter are generated as in eq. (3) with the replacement
@@ 2.1 Likelihood functions for the toy model
The maximum-likelihood estimator (MLE) for the likelihood parameters is the point $(\hat{\mu},\hat{{\theta}})$ that minimizes the negative log-likelihood function (NLL), which, apart from an irrelevant constant, can be written as
$-\ln L_{\rm P}(\mu,\theta)=\sum_{i=1}^{n}\left[f_{i}(\mu,\theta\,;\,t_{1i},t_{2i})-y_{i}\ln f_{i}(\mu,\theta\,;\,t_{1i},t_{2i})\right].$ (6)
Equation (6) neglects the fact that $t_{ji}$ are also random variables. A straightforward way to account for their stochastic nature is the Barlow-Beeston approach BARLOW1993219 , which consists in treating the vectors ${\bf t}_{j}=(t_{j1},\ldots,t_{jn})$ as independent Poisson-distributed random variables with unknown mean values $\boldsymbol{\theta}_{j}$, therefore extending the likelihood to include an extra contribution:
@@
A major drawback of eq. (7) is the growth in dimensionality: the number of likelihood parameters increases from $p=2$ to $p=2(n+1)$. However, as discussed in ref. BARLOW1993219 , numerical minimization of eq. (7) can be still recast as a two-dimensional optimization problem, at the cost of solving a non-linear system of equations BARLOW1993219 at each step of the minimization, which can be challenging numerically BBlite . For this reason, a simplification to eq. (7) has been proposed by Cowan BBlite , leading to the so-called Barlow-Beeston lite likelihood:
@@
$+\sum_{i=1}^{n}\beta_{i}m_{i}-m_{i}\ln\beta_{i},$ (8)
where the last term accounts for the total statistical uncertainty of the MC simulation in the $i^{\rm th}$ bin as represented by the effective number of MC events $m_{i}$ defined as
@@
2.1.1 Alternate likelihood parametrization
A data set distributed according to eq. (1) is clearly not sensitive to the signal normalization when $\epsilon=0$, which translates into a degeneracy of the likelihood function with respect to its parameters. For $\epsilon>0$, the problem is mathematically well defined, but the separation between the two processes becomes increasingly more difficult as $\epsilon$ approaches zero due to the large anti-correlation between $\mu_{1}$ and $\mu_{2}$. For example, for values of the $\epsilon$ parameter of ${\cal O}(1\%)$, the linear correlation $\rho$ between the two parameters is almost complete, e.g. $\rho=-0.99989$ for $\epsilon=0.03$. In fact, one might argue that the individual measurement of $\mu_{1}$ and $\mu_{2}$ is perhaps a poorly defined problem, given that the data is much more sensitive to their sum than to the difference. The role played by the choice of parameters is investigated by considering an alternate parametrization of the likelihood defined by the linear transformation

### v3

@@ Abstract
We consider the problem of setting confidence intervals on a parameter of interest from the maximum-likelihood fit of a physics model to a binned data set with a large number of bins, large event-counts per bin, and in the presence of systematic uncertainties modeled as nuisance parameters. We use the profile-likelihood ratio for statistical inference and focus on the case in which the model is determined from Monte Carlo simulated samples of finite size. We start by presenting a toy model in which the properties of widely used approximations of the profile-likelihood ratio in the asymptotic limit, which are commonly expected to hold in the high-statistics regime, are manifestly broken even if the numbers of events per bin in both the data and simulated samples are seemingly large enough to warrant their validity. We then move to the general setting to show how statistical uncertainties in the Monte Carlo predictions can affect the coverage of confidence intervals constructed in the asymptotic approximation always in the same direction, namely they lead to systematic under-coverage.
keywords:
HEP, Confidence Interval, Coverage, Barlow-Beeston.
journal: Nuclear Physics B
[…]
@@ 1 Introduction
Due to the complexity of modeling particle physics processes from generator up to detector-level, the function ${\bf f}(\mu,\boldsymbol{\theta})$ can rarely be computed analytically, leaving Monte Carlo (MC) techniques as the only viable way for its determination. Given that MC samples necessarily comprise a finite number of events, a MC-simulated model prediction will never be exact, thus introducing an additional source of systematic uncertainty. A general rule is that the size of the samples should exceed that of the data by at least a few times. However, such a request can turn out computationally expensive or even unfeasible for analyses performed on high-statistics samples of events, especially when CPU-intensive MC simulations are needed to achieve the desired level of theoretical accuracy. In such cases, one has to live with samples of MC-simulated events of size comparable or smaller than the data. The situation can become even more problematic when the simulated events are weighted [16].
@@
The correct frequentist approach to deal with the finite size of MC samples was first introduced by Barlow and Beeston in ref. [3]. In their seminal paper, the authors proposed the idea of treating the uncertain MC predictions as additional nuisance parameters of which the available MC sample represents an ancillary measurement. Since then, several improvements in the treatment of MC uncertainties have been proposed in the literature, with particular emphasis on the case of low-statistic samples and the proper handling of event weights [3, 5, 16, 1].
@@
In this paper, we focus on the problem of setting confidence intervals on a parameter of interest in the presence of nuisance parameters and in the case of high-statistics counting experiments, for which asymptotic formulas from the theory of maximum likelihood estimators [10, 28, 9] are expected to hold. We notice that the high-statistics case has received less attention in the literature compared to the low-statistics limit, where asymptotic formulas are more likely invalidated by low event counts. Furthermore, the high-statistics case is presently of great interest for precision measurements performed on large-scale data samples, such as those collected at the LHC, see e.g. refs. [14, 26] for a recent review on this subject.
@@
A relevant example of a precision measurement performed on a large-scale data sample, and with a large number of nuisance parameters, is represented by the measurement of the $W$ boson mass by the CMS Collaboration [26]: the CMS measurement analyzed a sample of LHC collision data consisting of about $10^{8}$ single-muon events, binned in the kinematics and charge of the muons, resulting in approximately $3\times 10^{3}$ bins, which are used to simultaneously measure the mass of the $W$ boson together with a few thousands of nuisance parameters via a maximum-likelihood fit. The model is constructed by using samples of MC-simulated events whose overall statistical power exceeds that of the data by a factor of about $4$. Alternate MC predictions corresponding to $\pm 1\sigma$ variations of each nuisance parameter are constructed by reweighting the nominal MC simulation. The effect of statistical fluctuations in the alternate templates used to construct the systematic variations was observed to lead to a moderate under-coverage of Hessian uncertainties, which required an inflation of the statistical uncertainty of the MC templates to restore the correct coverage [26]. The main scope of this paper is to study the origin of such an effect from a more general perspective.
@@
This paper is organized as follows. We start by presenting a hypothetical experiment which can be considered paradigmatic of our case of interest: the data set consists in a histogram with a large number of bins; the event counts in each bin are large; the physics model used to describe the data includes nuisance parameters and is determined from MC-simulated samples of size comparable to the data. We compute numerically the coverage of intervals constructed from CI-setting methods that rely on the asymptotic properties of maximum-likelihood estimators, such as those based on the Hessian matrix or the profile-likelihood ratio, and find that such intervals cover systematically less than their claimed confidence level. Then, we try alternative CI-setting methods that do not rely on asymptotic properties, such as those inspired by the unified approach by Feldman and Cousins [15] based on Neyman’s construction [21] and again find similar issues. Inspired by our findings, we propose a heuristic definition of a confidence interval that better approaches the correct Feldman-Cousins interval. In the second part of the paper, we study the same problem from a more general perspective to ascertain in which circumstances statistical fluctuations of the MC model can lead to a systematic under-coverage of asymptotic confidence intervals. We conclude by suggesting numerical tests to quantify the impact of MC statistical fluctuations for a generic problem, remarking that the validity of the asymptotic approximation implied by Wilks’ theorem should be carefully verified in the context of binned profile-likelihood analyses even when the number of measured events per bin is large and the statistical power of the simulated samples comparable to the data.
@@ 2 A paradigmatic toy model
Consider a hypothetical experiment measuring a set of event counts $y_{i}$, with $i=1,\ldots,n$, collected in the $n$ bins of a histogram. The measured events can be produced by either a signal or a background process, labeled by the indexes $j=1$ and $2$, respectively. The two processes are assumed to generate independent Poisson-distributed event counts $y_{ji}$ with expectation values in the $i^{\rm th}$ bin given by
$\langle y_{1i}\rangle\equiv\nu_{1i}=K,\;\;\;\langle y_{2i}\rangle\equiv\nu_{2i}=\begin{cases}(1+\epsilon)K&\;\;{\rm for}\;i<\frac{n}{2}\\ K&\;\;{\rm for}\;i\geq\frac{n}{2}\end{cases}$ (1)
where $|\epsilon|\ll 1$ and $K\gg 1$ are constant. For simplicity, we take $n$ as an even integer. In a real experiment, the constant $K$ would be related to e.g. beam flux, detector efficiency, cross sections, branching ratios, etc. In our study, we find it more convenient to express $K$ in terms of the total number of expected events ($N$) in the data sample, which can be easily determined in terms of $n$, $K$, and $\epsilon$ from eq. (1).
Let the purpose of this hypothetical experiment be a measurement of the unknown total number of signal events, henceforth denoted as parameter of interest (POI), with the unknown normalization of the background process playing the role of a nuisance parameter (NP). For this purpose, we will make use of a physics model which can predict the probability for a signal or background event to fall in a given bin. If this model could be derived directly from eq. (1), the data histogram could be re-binned in just two bins with no loss of information on either the POI or NP. However, the crucial assumption made here is that the physics model is not known analytically but can only be predicted from finite-size MC samples consisting of ${t}_{ji}$ simulated events in the $i^{\rm th}$ bin and for the $j^{\rm th}$ process. Since $t_{ji}$ are also random variables, they will differ from their true values by statistical fluctuations introduced by the MC generation, which could, in fact, justify preserving the full granularity of the data.
Given a set of MC predictions $t_{ji}$, the signal normalization can be extracted by parameterizing the unknown expectation value $\langle y_{i}\rangle$ in the $i^{\rm th}$ bin as:
@@
where strength modifiers $\mu_{1}$ and $\mu_{2}$ have been conveniently introduced for the two processes in place of their respective total number of events, and $k_{j}$ are constants that rescale the normalization of the MC samples to some reference value, which we are free to choose. In the following, we conveniently set $k_{j}$ equal to the ratio between the expectation value of $t_{ji}$ and its true value $\nu_{ji}$, so that by construction $f_{i}$ is an unbiased estimator of $\langle y_{i}\rangle$ for $\mu_{1}\equiv\mu_{1,{\rm t}}=1$ and $\mu_{2}\equiv\mu_{{2,\rm t}}=1$. With this choice, the constants $k_{j}$ represent the statistical power of the simulated samples relative to the data. For example, $k_{1}=k_{2}=1$ means that the total number of simulated events equals on average the number of data events, while values greater (smaller) than one indicate that more (less) events have been generated by the MC simulation than expected in data. Consistently with this assumption, numerical values for $y_{ji}$ and $t_{ji}$ in pseudo-experiments can be generated by drawing random integer numbers from the Poisson density with mean values $\nu_{ji}$ and $k_{j}\nu_{ji}$, respectively, that is:
${\rm Prob}\left[y_{ij}\right]=\frac{e^{-\nu_{ji}}\left(\nu_{ji}\right)^{y_{ji}}}{y_{ij}!}\;\;\;{\rm and}\;\;\;{\rm Prob}\left[t_{ij}\right]=\frac{e^{-k_{j}\nu_{ji}}\left(k_{j}\nu_{ji}\right)^{t_{ji}}}{t_{ij}!}.$ (3)
We notice that the rescaled MC predictions $T_{1i}=t_{1i}/k_{1}$ and $T_{2i}=t_{2i}/k_{2}$, commonly referred to as templates, also define the Jacobian matrix of the physics model ${\bf f}=(f_{1},\ldots,f_{n})$, with respect to its parameters $(\mu_{1},\mu_{2})$, that is
$\frac{\partial(f_{1},\ldots,f_{n})}{\partial(\mu_{1},\mu_{2})}=\begin{pmatrix}T_{11}&T_{21}\\ \vdots&\vdots\\ T_{1n}&T_{2n}\\ \end{pmatrix}$. (4)
The concept of Jacobian will be pivotal to section 3. Examples of templates $T_{ji}$ are shown in figure 1 for representative values of $N$, $\epsilon$, $n$, and $k_{j}$.
Figure 1: Examples of templates $T_{ji}$ generated as described in the text for $N=2\times 10^{6}$, $\epsilon=0.03$, $n=200$, and for two values of $k_{1}=k_{2}=k$, namely $k=1$ (left) and $k=10$ (right). The horizontal dashed lines in each plot provide the expected number of events per bin for the signal and background processes.
@@
In order to verify that the coverage of $I$ does not depend on $\mu_{\rm t}$, two additional independent data sets $\bf y$$_{\pm}$ are also considered. The latter are generated similarly to the left of eq. (3) with the replacement
@@ 2.1 Likelihood functions for the toy model
[…]
$\ln L_{\rm P}(\mu,\theta)=\sum_{i=1}^{n}\left[f_{i}(\mu,\theta\,;\,t_{1i},t_{2i})-y_{i}\ln f_{i}(\mu,\theta\,;\,t_{1i},t_{2i})\right].$ (7)
Equation (7) neglects the fact that $t_{ji}$ are also random variables. A straightforward way to account for their stochastic nature is the Barlow-Beeston approach [3], which consists in treating the vectors ${\bf t}_{j}=(t_{j1},\ldots,t_{jn})$ as independent Poisson-distributed random variables with unknown mean values $\boldsymbol{\theta}_{j}$, therefore extending the likelihood to include an extra contribution:
@@
A major drawback of eq. (8) is the growth in dimensionality: the number of likelihood parameters increases from $p=2$ to $p=2(n+1)$. However, as discussed in ref. [3], numerical minimization of eq. (8) can be still recast as a two-dimensional optimization problem, at the cost of solving a non-linear system of equations [3] at each step of the minimization, which can be challenging numerically [7]. For this reason, a simplification to eq. (8) has been proposed by Conway [7], leading to the so-called Barlow-Beeston lite likelihood:
@@
$+\sum_{i=1}^{n}m_{i}\left[\beta_{i}-\ln\beta_{i}\right],$ (9)
where the last term accounts for the total statistical uncertainty of the nominal MC simulation in the $i^{\rm th}$ bin as represented by the effective number of MC events $m_{i}$, defined as [^1] The effective number of MC events can be defined such that its relative uncertainty equals the expectation for $m_{i}$ unweighted events: $\frac{1}{\sqrt{m_{i}}}=\frac{\sqrt{\rm Var\left[T_{1i}+T_{21}\right]}}{T_{1i}+T_{2i}}=\frac{\sqrt{\frac{t_{1i}}{k_{1}^{2}}+\frac{t_{21}}{k_{2}^{2}}}}{\frac{t_{1i}}{k_{1}}+\frac{t_{21}}{k_{2}}}$ (10) For a MC sample consisting of $M$ weighted events, each one having weight $w_{k}$, Eq. (11) can be generalized to $m_{i}={\left(\sum_{k=1}^{M}w_{k}\right)^{2}}/\left({\sum_{k=1}^{M}w_{k}^{2}}\right)$.
@@
A summary of the bin-dependent variables introduced in the text is provided in Table 1.
Variable Definition Role
$\nu_{ji}^{(\pm)}$ True num. of events from the $j^{\rm th}$ process exact
$y_{i}$ Num. of data events random
$t_{ji}$ Num. of MC events from the $j^{\rm th}$ process random
$T_{ji}$ Template value for the $j^{\rm th}$ process ($t_{ji}/k_{j}$) random
$m_{i}$ Effective num. of MC events random
$\theta_{ji}$ Barlow-Beeston NP for the $j^{\rm th}$ process parameter
$\beta_{i}$ Barlow-Beeston lite NP parameter
$f_{i}$ Parametric model prediction for $\langle y_{i}\rangle$ exact/random
Table 1: Glossary of the variables defined for the $i^{\rm th}$ bin, as discussed in the text. In the last line, the model prediction $f_{i}$ should be considered as a fully parametric model (i.e. “exact”) for the Barlow-Beeston likelihood and “random” for the Barlow-Beeston lite.
2.2 Alternate likelihood parametrization
A data set distributed according to eq. (1) is clearly not sensitive to the signal normalization when $\epsilon=0$, which translates into a degeneracy of the likelihood function with respect to its parameters. For $\epsilon>0$, the problem is mathematically well defined, but the separation between the two processes becomes increasingly more difficult as $\epsilon$ approaches zero due to the large anti-correlation between $\mu_{1}$ and $\mu_{2}$. For example, for values of the $\epsilon$ parameter of ${\cal O}(1\%)$, the linear correlation $\rho$ between the two parameters is almost complete, e.g. $\rho=-0.99989$ for $\epsilon=0.03$. In fact, one might argue that the individual measurement of $\mu_{1}$ and $\mu_{2}$ is perhaps a poorly defined problem, given that the data is much more sensitive to their sum than to the difference.
The role played by the choice of parameters is investigated by considering an alternate parametrization of the likelihood defined by the linear transformation